OpenStudy (anonymous):

The curve y=x^2-3x-4 crosses the x-axis at P and Q. The tangents to the curve at P and Q meet at R. The normals to the curve at P and Q meet at S. Find the distance RS.

OpenStudy (phi):

I would find P and Q and sketch the graph just to get a feel for what is going on. can you find P and Q ? (factor the quadratic)

OpenStudy (anonymous):

(x-4)(x+1)

OpenStudy (phi):

yes. at the x-axis y=0 so we have (x-4)(x+1)=0 this means either x=4 or x=-1 this is a parabola, with a vertex at -b/2a with b= -3 and a=1 so it has a min at 3/2 y = (3/2)^2- 3*3/2 -4 = -6.25 Can you sketch out the problem?

OpenStudy (phi):